the accelaration of a particle is a=-4t if initially velocity is 8m/s the total distance covered by the particle till it comes to rest
 

Grade:11

1 Answers

Shailendra Kumar Sharma
6 years ago
a=-4t
dv/dt =-4t => v=-2t2+c @ t=0 v=8 that gives c=8 
so v= -2t2+8, this particle will be at rest i.e. v= 0 
t=2sec
So
now distacne ds/dt= v =-2t2+8
s=-2t3/3+8t+K
distance traveled before coming rest is s(2)-s(0) =-16/3+16+K -K =32/3 =10.66m